top of page
Date | Name | Affiliation | Other Observers | Instrument | Hours Observed | Hours Lost | Reason | Observing Program |
---|---|---|---|---|---|---|---|---|
09/19/2023 | Martha Leake | Valdosta State University | 8 students in Observational Technniques | CCD | 0 | 10 | Equipment Failure | Varied objects for 4101 class. Shutter problems required work in Peters Lab so the camera was brought down and the system powered down. We had clear skies in Valdosta! |
09/17/2023 | Robert Berrington | Ball State University | A486/586 Observing class | arc | 5 | 3 | Equipment Failure | Night was nice, but there appears to be a stuck shutter with the replacement camera. The shutter does open, but not all the way. The corners are frequently blocked by the camera shutter. We attempted to exercise the shutter, and tried different pointings, but the shutter continued to block the edges of the CCD image. |
09/14/2023 | William C Keel | University of Alabama | None | ARC CCD | 0 | 10 | Equipment Failure | ACE took camera components to Tucson for vacuum pumping and in hopes of repairing the connection issue. (Likewise dome circuit breaker for replacement). |
09/13/2023 | Julia Clark | ERAU/Otani/Pearson | N/A | CCD | 0 | 10 | Equipment Failure | CCD would “hang” on readout for all exposures despite multiple resets of ACE Node Service and reseting the rack a couple of times. |
09/10/2023 | Ken Rumstay | VSU (Emeritus) | None | CCD camera | 0 | 10 | Equipment Failure | Did not observe because of (a) poor weather (increasing clouds) and (b) connection issues with the Leach camera. |
09/08/2023 | Gary Henson | ETSU | none | Spectrograph | 1 | 6 | Clouds | Spectrograph controls all working normally now. Clouds from partly to threatening status most of night. Dome Controls not working-ACE notified. Leach CCD now exposing and reading out, but No DS9 display-ACE again notified. |
09/07/2023 | Walter Van Hamme | FIU | Samec | CCD | 0 | 10 | Software/Control PC Issues | Could not get camera to work for needed run. It worked earlier in the afternoon for biases and darks and failed when it was time to observe. Support could not get it up. Attention needed. |
09/06/2023 | Van Hamme | FIU | samec | CCD | 10 | 0 | No problems to report | Observed NSVS01363941 algol eclipsing binary. Covered Primary eclipse. Need several more nights to cover LC. |
09/02/2023 | William C Keel | University of Alabama | DragonCon Space Track | ARC CCD | 1 | 5 | Clouds | Reopened! Clear spot early in the night just long enough to verify pointing and find the BVRI focus is 58000. |
07/06/2023 | Walter Van Hamme | FIU | Samec | CCD | 5 | 4 | Clouds | Observing the short period Algol, NSVS01126493. |
06/26/2023 | Walter Van Hamme | FIU | Samec | CCD | 7 | 2 | Clouds | Observing the short period Algol, NSVS01126493. |
06/26/2023 | Walter VanHamme | FIU | Samec | ccd | 8 | 1 | Clouds | Observing the short period Algol, NSVS01126493. |
06/25/2023 | Walter Van Hamme | FIU | Samec | CCD | 9 | 0 | No problems to report | Observing the short period Algol, NSVS01126493.
|
07/07/2023 | William Keel | University of Alabama | 3 cats | ARC CCD | 5.5 | 3 | Software/Control PC Issues | Doing twilight flats, V exposure showed blank at level 140, aborted. Then "exposure start failed", so I cycled the telescope-outlet power and rebooted both computers. Still managed to set on main target by the end of twilight. During 60m exposure, the QSI guider silently disconnected (while leaving the display looking plausible) so I got 30" diagonal trails. (No screaming so as not to disturb the cats.) The second 60m exposure was successful with a disconnection event just after its end. ARC camera froze again around 0030 MST. I was unable to get it to complete an exposure and read out after that, even after the restarts that worked before. |
07/05/2023 | Gary Henson | ETSU | none | Spectrograph | 7.5 | 0 | No problems to report | Spectroscopy of WR stars, another excellent night for July. |
07/03/2023 | Gary Henson | ETSU | none | Spectrograph | 7 | 0 | No problems to report | Spectra of WR stars, a very good night, warm but clear and autoguider working well. |
07/02/2023 | Gary Henson | ETSU | none | Spectrograph | 3 | 5 | Clouds | Spectroscopy of WR stars, instruments working fine, but cloudy patches forming and moving in from Northeast middle of night. |
06/27/2023 | Robert Berrington | Ball State University | none | arc | 7 | 0 | No problems to report | Short Period Variable stars. |
06/25/2023 | Lindsay Spence | Embry-Riddle Aeronautical University | Tomomi Otani | CCD | 8 | 0 | No problems to report | Clear night |
06/24/2023 | Lindsay Spence | Embry-Riddle Aeronautical University | Tomomi Otani | CCD | 8 | 0 | No problems to report | Clear night |
06/22/2023 | William Keel | University of Alabama | 2 cats | ARC CCD | 7 | 0 | No problems to report | Windy, best results looking downwind to N. Guider x-stage labels are opposite the EW directions of buttons for motion; also, had to change X and Y guider speeds from 1,1 to -1, -1. RAdmin from telescope computer disconnected with annoying frequency. Focus 454000, guider focus 3000. |
06/20/2023 | S Johns | Texas A&M Commerce | Lena D, Shayaan Z, Dr. Williams | Andor | 7 | 0 | No problems to report | White dwarf relative photometry. |
06/18/2023 | Lindsay Spence | Embry-Riddle Aeronautical University | Julia Clark | CCD | 6 | 2 | Clouds | Lots of Clouds |
06/17/2023 | Lindsay Spence | Embry Riddle Aeronautical University | Julia Clark | 0.9 m | 8 | 0 | No problems to report | None |
06/14/2023 | Xianming Han | Butler University | Spencer Decker | CCD | 8 | 0 | No problems to report | eclipsing binary |
06/13/2023 | Xianming Han | Butler University | Spencer Decker | CCD | 8 | 0 | No problems to report | eclipsing binary |
06/12/2023 | Xianming Han | Butler University | Spencer Decker | CCD | 8 | 0 | No problems to report | eclipsing binaries |
06/11/2023 | Ken Rumstay | VSU (emeritus) | None | Leach CCD camera | 5 | 5 | Clouds | Tried to observe AGNs, despite increasing cloud cover. Gave up around 5:00 am EDT. The telescope appears to systematically point north of target after slewing, by about two arcminutes. |
06/07/2023 | Gary D Henson | ETSU | None | Spectrograph - 2nd Half | 4.5 | 0 | No problems to report | Spectra of Massive Stars - Poor seeing and Cirrus Clouds but FLI ccd now cool and spectrograph working nominally! |
06/07/2023 | Kaylee Grace | MCAO/UD | none | Leach | 8 | 0 | No problems to report | Observed pulsating white dwarf GD358. |
06/03/2023 | Kaylee Grace | MCAO | Judi Provencal, Millie Dill, Zach Savery, Tali Natan | Leach | 7 | 0 | No problems to report | Observed pulsating white dwarf GD358 |
06/02/2023 | Kaylee Grace | MCAO | Judi, Henry, Jarrod, Millie | Leach | 0 | 5.5 | Clouds | N/A |
06/01/2023 | Kaylee Grace | MCAO | Millie, Judi, Henry, Zach, Tali, Jarrod | Leach | 9.5 | 0 | No problems to report | Imaged pulsating white dwarf GD358 |
05/28/2023 | Todd Hillwig | Valparaiso University | Krusza, Blanton | ARC camera | 6 | 2 | Equipment Failure | Observing on-site. Mirror covers would not open (only one petal was operating). Notified ACE. Peter drove up and found it appeared to be a little bit of corrosion on one of the petal switches. Exercised the switch once and it started working again. Continued exercising to make sure it would keep working. No problems the rest of the night. |
05/27/2023 | Todd Hillwig | Valparaiso University | Lilly Blanton, Sami Krusza | ARC camera | 8 | 0 | No problems to report | M101 supernova and PN central stars. |
05/27/2023 | Tomomi Otani | Embry-Riddle Aeronautical University | Julia Clark, Lindsay Spence | CCD | 8 | 0 | No problems to report | Hot subdwarf time series |
05/27/2023 | Tomomi Otani & team | Embry- Riddle Aeronautical University | Julia Clark, Lindsay Spence | CCD | 8 | 0 | No problems to report | N/A |
05/26/2023 | Tomomi Otani | ERAU | Julia Clark, Lindsay Spence | CCD | 8 | 0 | No problems to report | Hot subdwarf time series |
05/26/2023 | TOMOMI | ERAU | Julia Clark, Lindsay Spence | CCD | 8 | 0 | No problems to report | Hot subdwarf time series |
05/24/2023 | James Webb | Florida International University | none | Leach CCD | 4 | 0 | No problems to report | Observing Program - Monitoring Blazars: PKS 1156+295, and 3C 279.
Problems: I couldn't figure out the syntax on how to automatically increment images number in the Leach software on the CCD computer. For the record the syntax is: imagename_{{seq:3}}.fts Turns out the .fts extension was the secret, thanks Brian and Bill.
The images came down with 1.4 (lower right hand quadrant missing). I fixed it by going into "full frame" and changing the "readout mode" to amplifier A instead of "quad readout", They seemed to come down complete and normal.
Brian said he would look at it this morning. |
05/18/2023 | William Keel | University of Alabama | 2 cats | ARC CCD | 0 | 9 | Clouds | Apparently this year's monsoon is trying a beta test. |
05/17/2023 | Gillam | Embry -Riddle Aeronotical University | Divyeshwari Vansadia, Aakasah Thiyagarajan | ARC Camera | 0 | 11 | Clouds | Microvaribales |
05/16/2023 | Stephen Gillam | ERAU | Aakash and Divyeshwari. | Andor Camera | 0 | 12 | Clouds | Micro-variable stars. |
05/15/2023 | Stephen Gillam | ERAU | Divyeshwari and Aakash | Andor Camera | 5.5 | 6.5 | Clouds | Micro-variable stars. |
05/14/2023 | Xianming Han | Butler Univ | none | ccd | 7 | 1 | Equipment Failure | filter wheel stuck once |
05/09/2023 | Matt A Wood | Texas A&M University-Commerce | NA | Leach | 2.0 | 8 | Wind | Time series. Wind hit 40 mph so I closed. Checked a couple times later in night and still high so didn't reopen as it was strong and from the south - didn't want to risk the mirror cover petals getting damaged. |
05/09/2023 | William C. Keel | University of Alabama | none | Andor CCD #2 | 8.5 | 0 | No problems to report | NGC 4945. Focus 129000, Without guider, 3-5 minute exposures at dec -45 have about 50% success rate in delivering round images using custom track rates. DS9 and MaximDL may have quit updating toward end of last sequence. |
05/08/2023 | Tom Rutherford | ETSU | None | CCD | 3 | 0 | No problems to report | Asteroid. |
05/07/2023 | Nayra Rodríguez | Instituto de Astrofisica de Canarias (IAC) | N/A | ARC CCD | 5 | 0 | No problems to report | Student research projects on exoplanet transit, 2nd half night. |
05/07/2023 | Matt Wood | Texas A&M University-Commerce | NA | Leach | 4 | 0 | No problems to report | Time series of white dwarf. This was a half night, so IAC observer took over at about 07:00 UT. Best focus early night was 448500 |
05/03/2023 | Nayra Rodríguez Eugenio | Instituto de Astrofisica de Canarias (IAC) | NA | ARC CCD | 0.5 | 4 | Humidity | Exoplanet transit, 2nd half night. Only observed from 7:30 to 8:10 UT. Domes closed due to high humidity and variable conditions. Focus 445000. |
05/01/2023 | Tom Rutherford | ETSU | None | CCD | 2 | 6 | Connection Issues | About two hours after starting, the CCD computer disconnected. I made several attempts to reconnect over the next hour or so but was unable to do so. Tried again a couple of times this morning (the next day) but still am unable to connect. Everything was working fine until this happened. |
04/28/2023 | William Keel | University of Alabama | N/A | ARC CCD | 6 | 0 | No problems to report | AGN with extended emission regions. Focus 445000, guide focus -18500. Filter wheel 1 needed INIT after almost every movement to put filters in position. |
04/27/2023 | Matt A Wood | Texas A&M University-Commerce | NA | Andor | 10 | 0 | No problems to report | Time series observations of white dwarf and cataclysmic variable. Focus 450000 at beginning of night, 445000 end of night. |
04/26/2023 | Matt A Wood | Texas A&M University-Commerce | NA | Andor | 10 | 0 | No problems to report | Time series observations of cataclysmic variables. Focus 444000 |
04/26/2023 | Matt A Wood | Texas A&M University-Commerce | NA | Andor | 10 | 0 | No problems to report | Time series observations of cataclysmic variables. Focus 443500. No problems. |
04/24/2023 | james webb | Florida International Universigy | none | Leach CCD | 4 | 0 | No problems to report | wasn't able to find guide star, guide stage wouldn't go to east positions, only west.
Wasn't able to save images from Telescope CCD control menu, had to use the camera menu on the ccd computer. Not sure what the issue was. But got the images I needed.
|
04/16/2023 | Xianming Han | Butler Univ | none | CCD | 8 | 1.5 | Equipment Failure | filter wheel keeps getting stuck when moving from Bessel I to Bessel B - emailed ACE about it. |
04/21/2023 | Todd Hillwig | Valparaiso University | Public demonstration | ARC camera | 3 | 0.5 | Other | Filter wheel did not seem to set properly a few times. Not a real problem since I was just doing a public demo. Quit early since Board meeting in the morning. |
04/20/2023 | William Keel | University of Alabama | 2 cats | ARC CCD | 7 | 0 | No problems to report | Galaxies with extended AGN-ionized clouds. Focus 399500, guider focus -16500. Focus shifted by 4000 units when moving telescope by 2 hours RA near midnight. Had to change guider setting X Speed from -1 to +1. |
04/19/2023 | Ken Rumstay | VSU (Emeritus) | None | CCD camera | 8 | 0 | No problems to report | AGN observations.
Everything worked find until the end, when I tried to observe 3C-390.3, which was near the NCP (HA ~ -4 hours, Dec. = +79 degrees). No stars appeared in the image. I suspect the dome may not have been aligned with the telescope, but it was pretty cloudy towards the north at that time. |
04/18/2023 | Kurtis Williams | TAMUC | Fangyi Zhu | ARC camera | 5 | 0 | No problems to report | White dwarf time series photometry; First training night for Zhu. No problems. |
04/17/2023 | Matt A Wood | Texas A&M University-Commerce | NA | CCD | 11 | 0 | Other | Time series observations of cataclysmic variables.
A good night and nice to be "back in the saddle".
The only issue was that getting a focus was tedious. I eventually ran to zero and reset the encoder, so my best focus at end of night at 50F/10C was 395000 through the "IR Blocking" filter. |
04/15/2023 | Xianming Han | Butler Univ | none | CCD | 10 | 0.2 | Equipment Failure | eclipsing binary
filter wheel stuck |
04/14/2023 | Xianming Han | Butler Univ | none | CCD | 10 | 0.2 | Equipment Failure | eclipsing binary
filter wheel stuck |
04/13/2023 | Xianming Han | Butler Univ | none | CCD | 9 | 1 | Equipment Failure | eclipsing binary
filter wheel stuck |
04/09/2023 | Stephen Gillam | ERAU | Jacob Romero, Rhiannon Hicks, Stephen Barnier, Avery Reynolds. | ARC Camera | 5 | 7 | Clouds | Micro-variables. At one point, the focus position changed on its own. I had command it to the saved position and stop this command in order to get the spurious motion to stop. |
04/07/2023 | Michael Fauerbach | FGCU | NA | CCD | 8 | 0 | No problems to report | if you change the filters make sure the wheel goes to the correct position. Had it twice go to a different position in the wheel. Else no issues to report. |
04/06/2023 | Amy Lovell | Agnes Scott College | Oralyn, Yazmin, Daniela, AST200L | ccd | 6 | 0 | No problems to report | cluster and eclipsing binary imaging for student projects, a few clouds early on... |
04/05/2023 | Tom Rutherford | ETSU | None | CCD | 8 | 0 | No problems to report | Exoplanet Transit |
03/31/2023 | Tomomi Otani | ERAU | None | CCD | 6 | 3 | Software/Control PC Issues | Hot subdwarf photometry time series. Both ACE and ACE connect suddenly closed 5 times (both closed simultaneously). They closed when I was not touching anything at least 2 times, but I noticed that they also close when I clicked the "StellaCam 3" tab on "Auxiliary Instruments".
The filter wheel did not work well and the camera position moved between two filters at one point. It was fixed when I initialized the filters. |
03/30/2023 | Amy Lovell | Agnes Scott | AST200L | CCD | .4 | 8 | Clouds | student projects, clouds rolled in early |
03/29/2023 | William Keel | University of Alabama | 2 cats | ARC CCD | 7.5 | 0 | No problems to report | Ionized clouds around galaxies. Focus 460000, guider focus 0. Large focus jump to 456400 when moving to 3 hours east. |
03/26/2023 | Nayra Rodríguez | IAC | none | ARC CCD | 4.5 | 0 | No problems to report | Student research projects on exoplanet transits, 2nd half night |
03/23/2023 | William Keel | University of Alabama | none | ARC CCD | 4.5 | 0 | No problems to report | Emission regions around galaxies. Did some cloud-hopping at start of night. Condensation on CCD window, evaporated over 2 hours with shutter open. Dome azimuth was about 30 degrees off until setting to home. Focus 458000, guider focus 0. |
03/23/2023 | Martha Leake | Valdosta.State University | Tools class | ARC | 0 | 10,3 | Humidity | Corrections to this morning's report. I was able to do biases and darks (without opening the dome, of course). Domes remained closed past sunrise. |
03/23/2023 | Martha Leake | Valdosta State University | Ten students in Tools class for 1.3 hours | ARC camera | 0 | 10.3 | Humidity | Domes closed due to high humidity fog. RH began to fall at 5 am EDT but not enough to open domes before 3 MST. Need to return by 10 am. Never opened; did biases and flats with no problems. |
03/21/2023 | Robert Berrington | Ball State University | none | arc | 0 | 8 | Clouds | mountain closed due to clouds. Did not open. |
03/20/2023 | William Keel | University of Alabama | N/A | ARC CCD | 0 | 10 | Clouds | The overcast layer was late compared to the forecast, but still arrived before sunset. |
03/19/2023 | Michael Fauerbach | FGCU | NA | CCD | 0 | 10 | Clouds | Domes are CLOSED due to: threatening clouds
|
03/15/2023 | Martha Leake | Valdosta State Unniversity | none | ARC camera | 0 | 9.7 | Clouds | Asteroid photometry. Cloudy, rain, expected showers and thunderstorm, more showers and 100 % RH rest of night. Did not access computers; did not open. |
03/11/2023 | Stephen D. Gillam | ERAU | Stephen Barnier | ARC Camera | 0 | 12 | Equipment Failure | Micro-variables.
The dome shutters would not open and the dome would not rotate. The upper and lower shutter status indicators, and the dome azimuth status were blank. We stopped and restarted the ACE node service and then power-cycled the telescope and camera computers, and power-cycled the ACE control cabinet. All to no avail.
We took 120s exposures at the zenith and got images that contained just the bias level. This confirmed that the dome was closed.
|
03/10/2023 | Gary Henson | ETSU | none | Spectrograph | 0 | 6 | Clouds | Cloud bands, variable humidity, sky conditions never consistent. |
03/09/2023 | Tom Rutherford | ETSU | None | CCD | 4 | 0 | No problems to report | Variable star. |
03/06/2023 | Gary Henson | ETSU | none | Spectrograph | 0 | 8 | Clouds | Cloudy all night, some rain in the morning. Mountain closed to observing all night. |
03/04/2023 | Amy Lovell | Agnes Scott College | AST200L students | CCD | 7 | 0 | No problems to report | student imaging projects, steady focus, great night |
03/03/2023 | Priya Gokuldass | Florida Tech | None | CCD | 3 | 6 | Clouds | ULGRB host galaxies observation. |
03/02/2023 | Vero Petit | University of Delaware | Judi Provencal, students | spectrograph | 0 | n/a | Humidity | Did not open, the humidity was high. We used the opportunity to show students how to use the spectrograph. |
03/01/2023 | Martha Leake | Valdosta State University | Tools class til 10 pm | CCD | 0 | 9 | Wind | Not opening tonight. Winds up to 66 mph, humidity at 100%, blowing snow in weather predictions and domes now closed because of threatening clouds. We will not attempt to open or connect. Thursday looks similar on KPNO/Sells predictions (20% chance of more snow showers before 9 am). Leaving campus (9:15 MST). |
02/28/2023 | Gary Henson | ETSU | Rutherford | Andor CCD | 0 | 8 | Clouds | Exoplanet transit light curve. CCD cooler functioning, no equipment problems but clouds increasing at sunset and didn't clear until last few hours of night after transit ended. |
02/28/2023 | Todd Vaccaro | Ball State University | none | Arc CCD | 0 | 10 | Equipment Failure | Wrote off the night when I never got a response about the cooler. The fix by Brian late that night did not get through in the thread of emails, so I was never aware in time that the cooler of the ccd had been turned back on. Only after Bob Berrington forwarded an email to my gmail account did I see that it was turned back on. |
02/26/2023 | Robert Berrington | Ball State University | none | arc | 0 | 12 | Equipment Failure | The CCD camera appears to have warmed to 13 C despite the set point for the cooler being set to -110 C. All breakers on the switches are set as expected, but the camera was showing a strong gradient in the bias frames that ranged from ~2000 at low row numbers to 65535 (saturated) at 1500+ row numbers. This of course rendered the camera useless.
I did not open because of the nonfunctional camera. |
02/24/2023 | Tomomi Otani | ERAU | Erin Abraham | CCD | 0 | 8 | Equipment Failure | Public Outreach.
We noticed that the camera cooler is not working. We reboot the systems but it did not help.
Weather is not good anyway so we ended the observation. |
02/23/2023 | Kurtis Williams | TAMUC | Fangyi Zhu | ARC | 0 | 11 | Snow | Didn't open dome due to ice |
02/20/2023 | William Keel | University of Alabama | 2 cats | ARC CCD | 0 | 10 | Clouds | I had in mind looking at clouds around distant AGN rather than quite so locally. |
02/19/2023 | Xianming Han | Butler Univ | none | CCD | 0 | 10 | Clouds | did not open |
02/18/2023 | Xianming Han | Butler Univ | Kaylee Tenbarge | CCD | 1 | 9 | Clouds | eclipsing binary |
02/17/2023 | Xianming Han | Butler Univ | Eric Reed | CCD | 0 | 10 | Clouds | none. never opened |
02/16/2023 | Xianming Han | Butler University | none | CCD | 0 | 10 | Other | Ice on dome. |
02/15/2023 | Martha Leake | Valdosta State University | Tools class early | Arc camera | 0 | 7 | Snow | "Domes are closed because of Ice on Domes." Did not open. Left at 3 :15 am EST when message above did not change. Skies were clear, but temperatures still cold. Planned asteroid photometry. |
02/13/2023 | Kurtis Williams | TAMUC | None | CCD | 0 | 11 | Clouds | Forecast for clouds and possible winters precip; never opened. |
02/12/2023 | Kurtis Williams | TAMUC | none | CCD | 0 | 11 | Clouds | Weather forecast is for increasing clouds with chance of rain, so staying closed out of caution. |
02/11/2023 | Martha Leake | Valdosta State University | none | ARC camera | 11.5 | 0.5 | Other | I had to leave for an appointment at 7:30 am Sunday. No problems tonight, seeing was good early on, but quite variable this morning. Observed asteroids for light curve determinations.G82 is the number for SARA KP telescope in the asteroid finder programs (from Lowell and other observatories and users). Thank you to Ken for leaving such a well organized site! |
02/10/2023 | Ken Rumstay | VSU (Emeritus) | None | CCD camera | 5.5 | 6 | Clouds | AGN observations. No equipment problems; closed up when cloud cover became too thick. |
02/08/2023 | Todd Hillwig | Valparaiso University | None | ARC camera | 0 | 10 | Connection Issues | Had local network issues that didn't get fixed until this morning, so was unable to connect and observe. |
02/07/2023 | Tom Rutherford | ETSU | Gary Henson | CCD | 5 | 0 | No problems to report | Comet, M 31, NGC 1245. No issues at all. |
02/05/2023 | Priya Gokuldass | Florida Tech | None | CCD | 2 | 6 | Clouds | ULGRB host galaxy observation. Cloudy! |
02/02/2023 | Amy Lovell | Agnes Scott | AST200L students | CCD | 6 | 0 | No problems to report | Student projects and finding the comet. Focus wandered quite a bit, ended at 463000 |
02/01/2023 | Todd Hillwig | Valparaiso University | None | Andor CCD | 0 | 10 | Humidity | No observing, high humidity and variable conditions. |
01/30/2023 | Gary Henson | ETSU | none | Spectrograph | 3 | 5 | Humidity | Spectroscopy of Massive Stars, Humidity hovering near closing limit early evening, finally inched up to limit closing all domes a few hours after twilight. |
01/25/2023 | james webb | Florida International universoty | none | CCD | 5 | 0 | No problems to report | Blazars, AO 0235+164, and 0716+71 |
01/23/2023 | William Keel | University of Alabama | 2 cats | ARC CCD | 0.5 | 12 | Clouds | Weather socked in - just had time for a set of sky flats after filters were cleaned. |
01/21/2023 | Martha A. Leake | Valdosta State University | none | Arc camera | 12,4 hours | 0 | No problems to report | Asteroid light curves and Comet E3 pictures. The KPNO domes were open, but temperatures near 32 F-34F; low humidity and clear skies. Productive night. Had to stop and restart slew to zenith park but no problems after that. The bright stars in my asteroid field left ghost streaks on the longer darks. Red message on my first (old) computer that Windows 10 was required; so I switched to Ken's computer. CCD Temperature sometimes displayed as 1000C, rather than -109 C; common glitch in display. A good night! |
01/21/2023 | Todd Vaccaro | Ball State University | N/A | ARC | 9 | 2 | Humidity | EBs; ACE was not working properly on either machine or up when I logged in, so both were rebooted and then seemed fine. The CCD pc had several pages of browser tabs with login failures, the usual SARA login/pw weren't working as they did on the tele pc. Tele seemed to have trouble slewing to first target, was hanging and had to be told twice. Weather station didn't seem to be working. |
01/19/2023 | Todd Vaccaro | Ball State University | N/A | ARC | 0 | 11 | Humidity | Eclipsing binary |
01/16/2023 | William Keel | University of Alabama | 2 cats | ARC CCD | 0 | 13 | Clouds | If I may quote the 4-meter operator's status for the night, "high humidity high winds dripping domes threatening clouds rain snow fog dangerous conditions". |
01/15/2023 | Priya Gokuldass | Florida Tech | None | CCD | 0 | 8 | Humidity | Ultra Long GRB host galaxies observation. 80% chances of rain and high humidity. |
01/14/2023 | Todd Vaccaro | Ball State University | none | ARC CCD | 11 | 0 | No problems to report | Eclipsing binary stars. Thin clouds most of the night and focus was off for R filter, ended with 473539 by end of the night, never better than 2.4". |
01/12/2023 | Gary Henson | ETSU | none | Spectrograph | 6 | 0 | Other | Test of spectrograph observations and performance after long shutdown due to fire. Details in saratech server email. Formal re-opening of observatory to regular operations after June 2022 fire. Noted debris over ccd dewar window and attempted to dislodge or at least describe behavior. |
01/10/2023 | William Keel | University of Alabama | 2 cats | ARC CCD | 2.5 | 10 | Clouds | Played cloud hopscotch for a test drive on long exposures. Focus 446500, guide focus 0. It's baaaaaack! |
12/17/2022 | Martha Leake | Valdosta.State University | none | Arc | 0 | 13 | Other | SARA 0.9-m still not in service because of the Contreras fire on the mountain. Conditions tonight look cloudy. Asteroid light curve observations will be tried when the telescope and computers and other facilities are operational. |
12/13/2022 | Ken Rumstay | VSU (Emeritus) | None | CCD camera | 0 | 12 | Other | Telescope still out of commission due to the Contreras fire in June. Even if that weren't the case. the Kitt Peak Site Weather Information website reports:
"Domes are CLOSED due to: high humidity, dripping domes, threatening clouds, ice on domes"
It's nice to see that website up and running again! |
10/07/2022 | Martha Leake | Valdosta State University | none | Arc | 0 | 10 | Other | Still not ready for observing due to Contreras fire damage to KPNO |
10/05/2022 | Priya Gokuldass | Florida Tech | None | CCD | 0 | 10 | Other | Observatory still closed due to Contreras fire. |
10/04/2022 | Ken Rumstay | VSU (Emeritus) | None | CCD camera | 0 | 11 | Other | KPNO still closed after Contreras fire in June. |
09/16/2022 | Ken Rumstay | VSU (Emeritus) | NOne | CCD canera | 0 | 10 | Other | Observatory still closed after the Contreras fire in June. |
07/08/2022 | Martha A. Leake | Valdosta. State University | none | Andor | 0 | 9 | Equipment Failure | Damage from Contreras fire; not connecting. |
07/02/2022 | Priya Gokuldass | Florida Tech | None | CCD | 0 | 10 | Other | Still closed due to Contreras fire |
06/29/2022 | Kurtis Williams | TAMUC | None | None | 0 | 8 | Other | Ongoing closure due to Contreras fire |
06/25/2022 | Kyle Pellegrin | Clemson University | N/A | N/A | 0 | 7 | Other | Observatory closed due to Contreras fire and power loss.
|
06/20/2022 | Robert Berrington | Ball State University | none | arc | 0 | 10 | Other | lost due to closure for mountain fire and power loss. |
06/19/2022 | Roebrt Berrington | Ball State University | none | arc | 0 | 10 | Other | lost due to closure for mountain fire and power loss. |
06/22/2022 | Kurtis Williams | TAMU-C | none | ARC | 0 | 9 | Other | Closure due to Contreras fire (night originally assigned to ERAU in exchange for my 6/23 night) |
06/21/2022 | Kurtis Williams | TAMU-C | none | ARC Camera | 0 | 9 | Other | Closure due to Contreras Fire |
06/13/2022 | Tom Rutherford | ETSU | None | CCD | 5 | 0 | No problems to report | Asteroid. |
06/12/2022 | Xianming Han | Butler Univ | Josh Wanninger | CCD | 8 | 0 | No problems to report | eclipsing binary |
06/09/2022 | Kyle Pellegrin | Clemson University | N/A | ARC CCD | 0 | 5 | Clouds | Clouds |
06/08/2022 | Robert Berrington | Ball State University | none | arc | 0 | 8 | Clouds | Short period variable star program.
Clouds did not open. |
06/07/2022 | Kurtis Williams | TAMU-C | REU Students | ACE CCD | 7.5 | 0 | No problems to report | White dwarf time series photometry. IR Blocking filter. Filter/Dewar had a hair or fiber that moved across field of view in three "jumps" in first hour of night - didn't affect our data - hurrah for point sources. Focus changed significantly over hour starting about 11pm MST - no sudden change in mountain weather at the time. Seeing was moderate, around 2.5" most of the night. |
06/04/2022 | Todd Hillwig | Valparaiso University | Scheel, Pavasars, Burbridge, Bakke | ARC CCD | 10 | 0 | No problems to report | PN Central Stars. Based on evening flats it appears that the dewar window has a LOT of dust. The images are pretty well saturated with donuts due to dust. |
06/04/2022 | Todd Hillwig | Valparaiso University | Will Bakke, Brook Burbridge, Emily Pavasars, Lane Scheel | ARC CCD | 5 | 4 | Other | Dust closure for first four hours. Rest of the night was clear, no problems. Observed PN central stars for variability. |
06/02/2022 | Xianming Han | Butler Univ | Josh Wanninger | CCD | 8 | 0 | Other | A line of shadow showed up since 2 nights ago (since night of June 1). It appears a wire (or something like that) fell onto the camera. |
06/01/2022 | Xianming Han | Butler Univ | Josh Wanninger | CCD | 7.5 | 0.5 | Equipment Failure | eclipsing binary
filter wheel stuck once |
05/31/2022 | Xianming Han | Butler Univ | Josh Wanninger | CCD | 8 | 0 | No problems to report | eclipsing binary |
05/30/2022 | Ken Rumstay | VSU | None | CCD camera | 9 | 0 | No problems to report | AGN program
One minor annoyance: I'm not sure why, but I was unable to change the sizes of any of the windows on the camera computer display. |
05/29/2022 | Robert Berrington | Ball State University | none | arc | 7 | 0 | Other | Short period variable stars
I was unsuccessful in getting the guider to successfully follow a star. In fact it appears to drive the star off the field rather than follow it. |
05/27/2022 | Martha Leake | Valdosta State University | none | ARC camera | 8.5 | 0 | No problems to report | Asteroids. Clouds at sunset continued then cleared somewhat. |
05/24/2022 | Todd Vaccaro | Ball State University | Nicolette Terracciano | ARC CCD | 8.5 | 0 | No problems to report | Eclipsing binaries. Thin clouds throughout the night and poor seeing with best focus around 2.5 arcsec. Forgot to turn autodome on for flats, but does that matter for zenith park? Dome was around 90 degrees and later when I closed I parked again at zenith and dome moved to 89ish, so I suppose no obstruction for earlier flats. |
05/20/2022 | Kyle Pellegrin | Clemson University | N/A | ARC CCD | 0 | 5 | Wind | Dome was not opened due to high winds and a report of dust/smoke in the air from the 4m telescope. |
05/14/2022 | Tom Rutherford | ETSU | None | CCD | 0 | 8 | Equipment Failure | Could not detect stars in images. Checked to see if the spectrograph was in the way-- it wasn't-- not sure at all what the issue was-- closed up after trying for a couple of hours. |
05/11/2022 | Tom Rutherford | ETSU | None | CCD | 4 | 3 | Clouds | Asteroids. Cirrus clouds at first but cleared after local midnight. No issues with the telescope or CCD. |
05/09/2022 | Kyle Pellegrin | Clemson University | N/A | ARC CCD & Spectrograph | 6.5 | 0 | No problems to report | Galaxy monitoring and spectroscopy practice. |
05/08/2022 | Roebrt Berrington | Ball State University | none | ARC | 7 | 0 | No problems to report | Short period variable stars |
05/05/2022 | Ken Rumstay | VSU (Emeritus) | Max the Cat | CCD camera | 8 | 1 | No problems to report | AGN Monitoring. First hour lost to stupidity. Observing went swimmingly once Todd reminded me to move the spectrograph out of the way. |
05/02/2022 | William Keel | University of Alabama | 3 cats | ARC CCD | 9 | 0 | No problems to report | Finished merger sample of Dead Quasar Survey. Oddity: TEL computer RAdmin connection froze often, and mostly when the CCD was reading out, requiring making a new connection. Focus 438200, guider focus -1600. |
05/01/2022 | William Keel | University of Alabama | 2 cats | ARC CCD | 2.5 | 6 | Other | Closed due to dust/smoke per 4m operator (Ken Rumstay pointed that out - at sunset a couple of domes including WIYN were open so I was lulled). Dead quasar survey. Focus=439500, guider focus 0. |
04/30/2022 | Stephen Gillam | ERAU | None | ARC Camera | 10 | 0 | No problems to report | rOAp stars.
Note: The CCD isn't quite lined up with the cardinal directions. N is to the right and W is up on images. There is a slight rotation of the CCD wrt the cardinal directions. |
04/24/2022 | Martha A. Leake | Valdosta State University | none | ARC camera | 9 | 1.5 | Other | Asteroid observations, no problems to report.
Another event causes the 1.5 h hours lost. Smoke was not an issue. |
04/23/2022 | Van Hamme | FIU | Samec | CCD | 10 | 0 | No problems to report | Observations of interacting binaries. GSC 30151656 and NSVS 10803881. |
04/21/2022 | Amy Lovell | Agnes Scott | Alex Yep, AST200L students | CCD + spectrograph demo | 4 | 4 | Wind | Training for AST 200L students with CCD, plus demo of spectrograph at the end
Closed due to high and increasing winds, but we got most of what we needed
ascom guide camera (for spectra) is not seeing the sky consistently, something odd is afoot, one time "move" made the star reappear, but it would disappear again with no explanation |
04/18/2022 | Kyle Pellegrin | Clemson University | N/A | ARC CCD | 4 | 0 | Clouds | Intermittent clouds, but was able to collect some good data. |
04/11/2022 | Ted von Hippel | ERAU | Leila Alamos, Yaneicy Jimenez | imaging | 6 | 0 | No problems to report | variable WD |
04/13/2022 | Stephen Gillam | ERAU | None | Andor Camera | 10 | 0 | No problems to report | rOAp stars. |
04/10/2022 | Tom Rutherford | ETSU | None | CCD | 4 | 0 | No problems to report | Asteroid. No issues. |
04/09/2022 | Walter VanHamme | FIU | Samec | CCD | 10 | 0 | No problems to report | Observations of interacting binaries. |
04/08/2022 | Ted von Hippel | ERAU | Lelia Alamos, Yaneicy Jimenez | imaging | 7 | 0 | No problems to report | Variable white dwarf photometry. The remainder of the night went to Kyle Pellegrin, a PhD student at Clemson for GRB work. |
04/07/2022 | Kyle Pellegrin | Clemson University | N/A | ARC CCD | 3.25 | 0 | Software/Control PC Issues | Had to restart ACE Node service to allow camera to download images to MaximDL. Other than that, everything went well. |
04/06/2022 | Kyle Pellegrin | Clemson University | N/A | ARC CCD | 8 | 0 | No problems to report | Monitoring galaxies for variability; exploring drizzle and lucky imaging. |
04/05/2022 | William Keel | University of Alabama | 3 cats | ARC CCD | 9 | 0 | No problems to report | Finished [O III] images for dead quasar survey merger phase. After moving to my third target and taking a test exposure, I watched as the focus readout shifted spontaneously by +35000 units. The actual focus did not change. Testing binning 2x2 during guiding worked well, speeding up the start of guiding. One long exposure wiped out by fast RA oscillations near 345W +41. |
04/04/2022 | William Keel | University of Alabama | 3 cats | ARC CCD | 9 | 0 | No problems to report | Dead quasar survey. Focus 395500, guider focus 0. |
04/03/2022 | Keefe Kamp | Florida Tech | Gokuldass, Tobolski | CCD/Spec | 2 | 8 | Clouds | Be Spectra and Tess Photometric follow-up planned however clouds appeared to come in because the CCD image began to look smeared and darker closed up after that |
04/02/2022 | Zorayda Martinez | Texas A&M University-Commerce | Kurtis Williams | Andor | 7 | 0 | No problems to report | white dwarf time series |
04/01/2022 | Zorayda Martinez | Texas A&M University-Commerce | Kurtis Williams | Andor | 7 | 0.5 | Other | white dwarf time series, time lost due to forgetting to click autosave |
03/25/2022 | XIANMING HAN | Butler University | none | CCD | 80 | 1.5 | Equipment Failure | filter wheel stuck |
03/31/2022 | William Keel | University of Alabama | 3 cats | ARCC CD | 9.5 | 0 | No problems to report | Dead quasar survey, When connecting in afternoon, connector client would not start. Brian Brondel had a look and found something messed up in the Windows registry. CCD focus 394000, guide focus=0. |
03/27/2022 | William Keel | University of Alabama | 3 cats | ARCCCD | 8.5 | 1 | Clouds | Dead quasar survey. Focus 394000. Guide focus -18000; had to change X rate in guide settings from 1 to -1, and reconnect QSI guide canera from MaxIm. |
03/23/2022 | Todd Vaccaro | Ball State University | Nicolette Terracciano | ARC CCD | 7.5 | .5 | Software/Control PC Issues | Eclipsing binary. We had some problem getting our target at a low altitude, possibly too low for dropout, but after homing the dome and moving the scope to confirm it was looking through the slit and going back to the target things were fine except for the beginning of a Bessel R and I sequence that got stuck trying to move the filter wheel to I, so I INIT wheel 1 after pausing the sequence and then restarting; things went smoothly for the rest of the night. |
03/21/2022 | Nayra Rodríguez Eugenio | IAC | none | none | 0 | 5 | Clouds | Clouds and humidity > 80% for the entire second half of the night. I didn't even connect to the telescope |
03/19/2022 | Tom Rutherford | ETSU | None | CCD | 4 | 6 | Clouds | Asteroid. High cirrus clouds-- sky became washed out once the moon came up. |
03/11/2022 | Ted von Hippel | ERAU | Leila Alamos, Yaneicy Jimenez, Kathryn Closius | imaging | 9 | 0 | No problems to report | Student programs on variable white dwarf and reflection nebula. |
03/12/2022 | Martha Leake | Valdosta State University | Logan Spooner | ARC camera | 9 | 1 | Other | No problems to report. Much time spent finding best focus (several times). No slips in focus noted. Needed to close up for 7:30 am commitment. Will transmit data shortly. Also, will stow telescope in 0 HA and -30 degree Dec position. |
03/08/2022 | Todd Hillwig | Valparaiso University | None. | ARC camera | 9.5 | 0.5 | Other | PN central stars. A few times during the night the tracking could not keep up with the object, creating E-W elongation/trails. This happened mostly in the west, but at intermediate altitudes in the NW. But also happened in the NE with one target. |
03/06/2022 | Ted von Hippel | ERAU | Leila Alamos, Yaneicy Jimenez | imaging | 6 | 0 | No problems to report | One filter reinit required. Some pedagogically useful operator errors from the students. Needed to move the spectrograph out of the way. |
03/07/2022 | Nayra Rodriguez Eugenio | IAC | none | ARC CCD | 4.5 | 0 | No problems to report | Student research projects on exoplanet transits, 2nd half night. |
03/04/2022 | Todd Vaccaro | Ball State University | Nicolette Terracciano, Jonathon Smith | ARC CCD | 2 | 9 | Wind | EB time of minimum, missed b/c we had to close due to high winds (lost 5 hrs and no flats at the end of the night), lost 3 hours at the beginning of night due to threatening clouds (no early flats), lost 1 hr b/c we thought clouds an issue, but the dome was not aligned with the scope (had to home dome, park scope, restart auto dome). Would be nice to have web cams work, but they don't appear to be working. We also had to login to each machine to start and nothing was up and running, which was confusing and couldn't get the telescope machine to initiate ACE at first, so that machine was rebooted. The night will be a complete loss if we don't have someone offer some Bessel R and I flats for us to use. |
02/28/2022 | William Keel | University of Alabama | 3 cats | ARC CCD | 11 | 0 | No problems to report | Dead quasar survey. Spent an hour chasing focus, discovered it had changed by 110,000 units since my last KP night, then had a couple of hours of the worst seeing I've ever experienced at Kitt Peak. Focus eventually settled to 39250, slid to 385200 later in the night (kind of like airport gates in Airplane). Guide stage still seems stuck at edge of field; focus was -17600. With fixed guider position, trick of binning guide exposures 2x2 in Setup panel (I think I heard that from Kurtis Williams) really helped for faint guide tars. |
02/23/2022 | Ken Rumstay | VSU | None | CCD camera | 0 | 11 | Humidity | AGN's |
02/22/2022 | Gary Henson | ETSU | None | Spectrograph | 0 | 9 | Wind | High humidity off and on but winds at 4m up to 60 mph |
02/20/2022 | Gary Henson | ETSU | Rutherford | CCD and Spectrograph | 0 | 9 | Clouds | Threatening clouds early, clear for a couple hours but more clouds moving in soon after that. Not enough good weather to open. |
02/19/2022 | Dr. Alexandra Yep | Agnes Scott College | none | Spectrograph | 4 | 0 | Software/Control PC Issues | I was trying to do spectroscopy and could not figure out how to get a decent spectrum. I don't know what focus to use or what I'm doing wrong. Besides that though, I think I maybe needed the StellaCam3? Every time I clicked that tab, the whole Ace program crashed. That might be a problem? |
02/17/2022 | Ted von Hippel | ERAU | Elvis Santana | imaging | 6.5 | 0 | No problems to report | Observation (Seyfert type 2 galaxies and quasars) went well. SARA-KP was responsive, imaging went off without a kink, and weather conditions were good. |
02/18/2022 | Martha Leake | Valdosta State University | none | Andor | 11.2 | 0 | No problems to report | Asteroid light curve work for three objects and images for a Stars and Galaxies class. All went well but needed to connect to camera computer near start up; used the Ace connector from the start button. Pointing slightly off center. Stowed the telescope. |
02/16/2022 | Xianming Han | Butler University | Josh Wanninger | CCD | 0 | 10 | Humidity | did not open |
02/15/2022 | Xianming Han | Butler University | Kaylee TenBarge | CCD | 4 | 6 | Wind | eclipsing binaries |
02/14/2022 | Xianming Han | Butler University | Kira Baasch | CCD | 10 | 0 | No problems to report | eclipsing binaries |
02/13/2022 | Xianming Han | Butler University | Josh Wanninger | CCD | 10 | 0 | No problems to report | eclipsing binaries |
02/12/2022 | Xianming Han | Butler University | Maya Leonard | CCD | 10 | 0 | No problems to report | eclipsing binaries |
02/11/2022 | Xianming Han | Butler University | Kaylee TenBarge | CCD | 10 | 0 | No problems to report | eclipsing binaries |
02/10/2022 | Xianming Han | Butler University | Kira Baasch | CCD | 10 | 0 | No problems to report | eclipsing binaries |
02/09/2022 | Xianming Han | Butler University | Josh Wanninger | CCD | 10 | 0 | No problems to report | eclipsing binaries |
bottom of page